In a Wheatstone’s bridge, there resistances P, Q and R connected in the three arms and the fourth arm is formed by two resistances S1 and S2 connected in parallel. The condition for bridge to be balanced will be
?

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
In a Wheatstone’s bridge, there resistances P, Q and R connected in the three arms and the fourth arm is formed by two resistances S1 and S2 connected in parallel. The condition for bridge to be balanced will be
given by\frac{P}{Q}=\frac{R*S1S2}{S1+S2}
Thanks and Regards
Shaik Aasif
askIITians faculty
SHAIK AASIF AHAMED
9 years ago
Hello student,
In a Wheatstone’s bridge, there resistances P, Q and R connected in the three arms and the fourth arm is formed by two resistances S1 and S2 connected in parallel. The condition for bridge to be balanced will be
given by\frac{P}{Q}=\frac{R*(S1+S2)}{S1S2}
Thanks and Regards
Shaik Aasif
askIITians faculty